# Trust, Because You Can’t Verify: Privacy and Security Hurdles in Education Technology Acquisition Practices

### v1

@@ A.3 Post-Acquisition Phase
Table 1: Study Participant Demographics
ID Current Role Role in EdTech Acquisition Years of Exp. No. of EdTech Handled Region
P1 Chief Technology Officer (CTO) Oversees security review 23 100 West Coast
P2 Chief Information Officer (CIO) and VP for Information Technology Oversees security audits 15 Multi 100s East Coast
P3 Deputy Chief Information Officer (Deputy CIO) Oversees security audits 23 Mid-west
P4 Chief Information Security Officer (CISO) Oversees security reviews and audits 3 Multi 100s South
P5 Chief Information Security Officer (CISO) Reviews technologies in procurement that handle sensitive data 10 Mid-west
P6 Deputy Chief Information Officer (Deputy CIO) Implements privacy controls in contracts 4 Multi 100s South
P7 Senior Director of Learning Oversees procurement process 7 Multi 100s South
P8 Senior Director of Learning Experience and IT Services Oversees procurement 4 Multi 100s South
P9 Executive Director of Data Analysis Handles on premise technology data storage 22 Multi 100s South
P10 Enterprise Partner Part of negotiations with vendors 13 Multi 100s South
P11 Deputy Chief Information Officer (Deputy CIO) Oversees security reviews and purchasing committees 25 100 Mid-west
P12 Chief Information Security Officer (CISO) Oversees security reviews and data handling 12 Multi 100s North Eastern
P13 IT Support Implements edtechs after acquisition 5 100 South
@@ Appendix B Supplementary Study Materials
The codebook for this study, containing theme names, category names, category descriptions, and open codes, can be accessed here: https://osf.io/84y2h/?view_only=1f0f36c5673a4ebdb78e49e22f319b66

### v2

@@ A.3 Post-Acquisition Phase
Table 1: Participant Demographics
ID Current Job Title Role in EdTech Acquisition (Years in current Role) Years of Exp. in EdTech EdTech Handled Region
P1 Chief Technology Officer (CTO) Oversees security review (4) 23 ~100 West Coast
P2 Chief Information Officer and VP for Information Technology Oversees security audits (4) 15 200+ East Coast
P3 Deputy Chief Information Officer Oversees security audits (2) 23 - Mid-west
P4 Chief Information Security Officer Oversees security reviews and audits (3) 3 200+ South
P5 Chief Information Security Officer Oversees procurement of EdTech handling sensitive data (5) 10 - Mid-west
P6 Deputy Chief Information Officer Implements privacy controls in contracts (4) 4 200+ South
P7 Senior Director of Learning Oversees procurement (7) 7 200+ South
P8 Senior Director of Learning Experience and IT Services Oversees procurement (4) 4 200+ South
P9 Executive Director of Data Analysis Handles on premise technology data storage (5) 22 200+ South
P10 Enterprise Partner Negotiations with vendors (13) 13 200+ South
P11 Deputy Chief Information Officer Oversees security reviews and purchasing committees (9) 25 ~100 Mid-west
P12 Chief Information Security Officer Oversees security reviews and data handling (2) 12 200+ North-east
P13 IT Support Handles Edtech post-acquisition (5) 5 ~100 South
@@ Appendix B Supplementary Study Materials
The codebook for this study, containing theme names, category names, category descriptions, and open codes, along with the recruitment email can be accessed here: https://osf.io/84y2h/?view_only=1f0f36c5673a4ebdb78e49e22f319b66